# Optimal Transmitter Placement in Realistic Urban Environments

### 30 Apr 2026 v1

@@ I-A Relation to Prior Work
Physically grounded optimization is increasingly enabled by digital twins and ray-tracing frameworks like Sionna RT [17], which incorporate site-specific material properties and antenna orientations [2, 44]. Such high-fidelity modeling is critical in mmWave and reconfigurable intelligent surface (RIS) networks [42, 33, 13, 4].
Methods that utilize expensive gradient based searches have been explored to optimize transmitter locations [3]. However, these methods are myopic and require a-priori knowledge of the required number of transmitters. Our method scales by approximating interference through the physical overlap of ray-traced areas, enabling a high-fidelity optimization loop at a fraction of the computational cost.
@@ I-B Contributions
Realistic Validation: The approach is validated using the Sionna RT ray-tracing framework applied to realistic 3D models of San Francisco and Florence, demonstrating significant throughput improvements over existing deployments by AT&T and T-Mobile in SF, and Iliad in Florence. For example, IA-SPA achieves mean throughput gains of up to 70% in dense urban scenarios and edge-rate improvements exceeding 130%, while simultaneously keeping interference low.
@@ II-A Propagation Model
We define the propagation function $P(y,x)$, which represents the signal power received at location $x\in\Omega$ when the signal originates from a source at location $y\in\Omega$ and propagates through the environment $\Omega$ according to the material-dependent properties $\mu(x)$.
@@
In the following, $P(y,x)$ is considered to be the Signal-to-Noise Ratio (SNR) at location $y$ from a transmitter at $x$. Without loss of generality, we assume a normalized noise power of $\sigma^{2}=1$, such that $P(y,x)$ also corresponds to the received signal power at $y$. Fig. 2 illustrates $P(y,x)$, computed for the environment in Fig. 1 using Sionna. Each panel shows the propagation for a different transmitter location y (red marker). The resulting color map at x represents the data rate, demonstrating how environmental geometry and material properties transform a simple source into a complex coverage field.
Figure 2: $P(y,x)$ for an urban environment. The red marker indicates the transmitter position $y$ and the colors represent the data rate at each location $x$ which corresponds to the values of $P(y,x)$.
@@ II-B Network Quality and SINR
To describe the interaction of multiple transmitters in the set $T=\{t_{1},t_{2},...,t_{n}\}$ in the environment, we use the function $\mathcal{P}(x,T)$ which describes the cumulative power at location $x\in\Omega$. In practice this is done by
@@ II-D Optimization Problem Formulation
If $\bar{W}$ is concave and for any $x\in\Omega$ $T\mapsto\mathcal{P}(x,T)$ is convex, then $S(T)$ is concave, and therefore the problems described by equations (11) and (12) are convex.
@@ IV-C San Francisco using $\mathcal{P}_{\text{MAX}}$
The IA-SPA optimization framework yields significant performance gains over existing infrastructure, nearly doubling mean throughput in established carrier scenarios. As detailed in Table II, the mean data rate increased by 70% in the AT&T scenario and by over 215% in the T-Mobile scenario. Even more striking is the improvement at the network edge (5th percentile), which saw increases of 134% and 758%, respectively. Remarkably, these capacity gains were achieved while simultaneously reducing network interference slightly by 2% for AT&T and a drastic 93% for T-Mobile, proving that the algorithm successfully disentangles overlapping signal footprints by exploiting site-specific 3D geometry.
@@
The IA-SPA framework delivers transformative capacity gains, nearly doubling mean throughput while simultaneously lowering the network interference floor. As detailed in Table II, the mean data rate increased by 70% for the AT&T scenario and over 215% for T-Mobile. These improvements are most pronounced at the network edge (5th percentile), where rates surged by 134% and 758%, respectively. Notably, these gains do not come at the cost of signal quality; interference decreased by 2% in the AT&T case and a drastic 93% in the T-Mobile case, proving that the algorithm successfully disentangles overlapping signal footprints.
@@ IV-E Florence with Exclusionary Zone
IV-F Florence with fixed Initial Deployment
In practice transmitters are rarely deployed into an empty environment with no previous transmitters present. More commonly, operators must perform incremental deployment, where new transmitters are positioned to augment a pre-existing configuration. To evaluate the IA-SPA’s performance in this context, we initialize the Florence environment with a fixed set of three transmitters (see Fig. 13), denoted as $T_{\text{fixed}}$, representing an existing baseline deployment. The algorithm is then tasked with selecting augment this configuration to maximize the aggregated network quality $\mathcal{S}(T\cup T_{\text{fixed}})$ using $\mathcal{P}_{\text{MAX}}$. For this numerical example we again use the Florence environment and the corresponding Iliad data set.
Figure 13: Top-down view of the Florence environment with an initial deployment of three fixed transmitters (indicated by red dots). These sites represent pre-existing infrastructure that the IA-SPA must account for when optimizing additional tower placements.
The IA-SPA framework effectively densifies existing networks while simultaneously lowering the interference floor—a critical requirement for modern urban deployments. When tasked with augmenting a fixed three-tower baseline ($T_{\text{fixed}}$), the algorithm successfully identified coverage gaps to provide a 54% improvement in the edge data rate (Table V). Rather than creating redundant signal overlap, the IA-SPA identified sites with high marginal utility, resulting in a 79% reduction in peak interference.
TABLE V: Performance Metrics: Fixed Baseline vs. IA-SPA Augmented Network (Florence)
Metric Reference IA-SPA Optimized Change
@@ IV-F Florence with fixed Initial Deployment
Mean Data Rate 39.15 50.53 $\uparrow$29.07%
Standard Deviation 29.23 29.20 $\downarrow$0.10%
Maximum Data Rate 204.78 209.03 $\uparrow$2.08%
Edge Rate (5%-tile) 7.43 11.51 $\uparrow$54.94%
@@
Mean Interference 8.81 6.16 $\downarrow$30.12%
Standard Deviation 18.24 6.67 $\downarrow$63.42%
Maximum Interference 506.44 104.86 $\downarrow$79.29%
Figure 14: Throughput distribution for the incremental deployment in Florence. The columns illustrate the coverage from the fixed initial towers (left), the final augmented network (center), and the absolute gain provided by the new sites (right).
Figure 15: Interference distribution for the incremental deployment. The IA-SPA identifies sites that minimize overlap with the fixed infrastructure, resulting in a significantly reduced interference standard deviation.
Figures 14 and 15 highlight this synergy. By accounting for existing signal footprints, the augmented network eliminates “voids” in the Iliad dataset, demonstrating that the algorithm is as capable of refining existing infrastructure as it is of designing greenfield deployments.
@@ IV-G Key Takeaways from Results
The most significant results are found in the San Francisco AT&T scenario. Unlike other baselines that exhibit massive coverage gaps, the AT&T deployment is dense and well-structured. Despite this high-quality starting point, the IA-SPA achieved a 70% increase in mean throughput and a 134% improvement in the edge (5th percentile) data rate. These gains are particularly convincing because they were achieved while simultaneously slightly reducing the interference (by 2%). This proves that the algorithm is performing a high-precision optimization that disentangles overlapping signal footprints by exploiting the site-specific 3D geometry of the city.
The comparison between the two accumulation models highlights the distinction between raw signal capacity and interference-limited performance. While the $\mathcal{P}_{\text{SUM}}$ configuration provides an idealized upper bound for throughput by assuming perfect coordination, it lacks the implicit pressure to avoid overlapping coverage areas. This is reflected in the numerical results. While IA-SPA still achieves significant gains when using $\mathcal{P}_{\text{SUM}}$, the improvements are noticeably lower than those seen under $\mathcal{P}_{\text{MAX}}$ (e.g. 43% mean rate increase vs 70% for AT&T). This discrepancy suggests that by optimizing for aggregate power rather than strongest serving link, the algorithm may place transmitters in locations that provide redundant coverage, rather than expanding the network’s effective footprint. Consequently. the $\mathcal{P}_{\text{MAX}}$ objective not only provides a more realistic representation of signal behavior, but also serves as a more effective driver for identifying strategic sites that mitigate the specific geometric ”dead zone” characteristic for dense urban topologies.
The Florence case studies serve to demonstrate that the results of IA-SPA remain robust under various operational constraints. When considering an exclusionary zone, the algorithm achieved an 55% edge-rate improvement under placement constraints. Similarly we also showed that the algorithm is capable of incremental deployment by computing optimal transmitters to be added to an existing infrastructure, which resulted in a 29% improvement in data rate and a 79% reduction in peak interference.

### 07 May 2026 v2

@@ I-A Relation to Prior Work
Physically grounded optimization is increasingly enabled by digital twins and ray-tracing frameworks like Sionna RT [17], which incorporate site-specific material properties and antenna orientations [2, 44]. Such high-fidelity modeling is critical in mmWave and reconfigurable intelligent surface (RIS) networks [42, 33, 13, 4]. Methods that utilize expensive gradient based searches have been explored to optimize transmitter locations [3]. However, these methods are myopic and require a-priori knowledge of the required number of transmitters. Our method scales by approximating interference through the physical overlap of ray-traced areas, enabling a high-fidelity optimization loop at a fraction of the computational cost.
@@ I-B Contributions
Realistic Validation: The approach is validated using the Sionna RT ray-tracing framework applied to realistic 3D models of San Francisco and Florence, demonstrating significant throughput improvements over existing deployments by AT&T and T-Mobile in SF, and Iliad in Florence. For example, IA-SPA achieves mean throughput gains of up to 70% in dense urban scenarios and edge-rate improvements exceeding 88%, while simultaneously keeping interference low.
@@ II-A Propagation Model
We define the propagation function $P(y,x)$, which represents the signal power received at location $y\in\Omega$ when the signal originates from a source at location $x\in\Omega$ and propagates through the environment $\Omega$ according to the material-dependent properties $\mu(x)$.
@@
In the following, $P(y,x)$ is considered to be the Signal-to-Noise Ratio (SNR) at location $y$ from a transmitter at $x$. Without loss of generality, we assume a normalized noise power of $\sigma^{2}=1$, such that $P(y,x)$ also corresponds to the received signal power at $y$. Fig. 2 illustrates $P(y,x)$, computed for the environment in Fig. 1 using Sionna. Each panel shows the propagation for a different transmitter location $x$ (red marker). The resulting color map at $y$ represents the data rate, demonstrating how environmental geometry and material properties transform a simple source into a complex coverage field.
Figure 2: $P(y,x)$ for an urban environment. The red marker indicates the transmitter position $x$ and the colors represent the data rate at each location $y$ which corresponds to the values of $P(y,x)$.
@@ II-B Network Quality and SINR
To describe the interaction of multiple transmitters in the set $T=\{t_{1},t_{2},...,t_{n}\}$ in the environment, we use the function $\mathcal{P}(y,T)$ which describes the cumulative power at location $y\in\Omega$. In practice this is done by
@@ II-D Optimization Problem Formulation
If $\bar{W}$ is concave and for any $y\in\Omega$ $T\mapsto\mathcal{P}(y,T)$ is convex, then $S(T)$ is concave, and therefore the problems described by equations (11) and (12) are convex.
@@ IV-C San Francisco using $\mathcal{P}_{\text{MAX}}$
The IA-SPA optimization framework yields significant performance gains over existing infrastructure, nearly doubling mean throughput in established carrier scenarios. As detailed in Table II, the mean data rate increased by 70% in the AT&T scenario and by over 215% in the T-Mobile scenario. Even more striking is the improvement at the network edge (5th percentile), which saw increases of 88% and 758%, respectively. Remarkably, these capacity gains were achieved while simultaneously reducing network interference slightly by 2% for AT&T and a drastic 93% for T-Mobile, proving that the algorithm successfully disentangles overlapping signal footprints by exploiting site-specific 3D geometry.
@@
The IA-SPA framework delivers transformative capacity gains, nearly doubling mean throughput while simultaneously lowering the network interference floor. As detailed in Table II, the mean data rate increased by 70% for the AT&T scenario and over 215% for T-Mobile. These improvements are most pronounced at the network edge (5th percentile), where rates surged by 88% and 758%, respectively. Notably, these gains do not come at the cost of signal quality; interference decreased by 2% in the AT&T case and a drastic 93% in the T-Mobile case, proving that the algorithm successfully disentangles overlapping signal footprints.
@@ IV-E Florence with Exclusionary Zone
IV-F San Francisco with fixed Initial Deployment
In practice, transmitters are rarely deployed in empty environments without existing infrastructure. Instead, operators typically face a continuous deployment setting, where new transmitters must be placed to augment an existing network. To evaluate IA-SPA in this more realistic scenario, we initialize the system with the 9 transmitter locations deployed in San Francisco by AT&T (shown in Fig. 6). Starting from this baseline, IA-SPA sequentially determines up to 5 additional transmitter locations to improve network coverage. We assess performance by comparing against random transmitter placement, reporting results averaged over 10 independent trials to reduce variability due to randomness.
Figure 13: Average data rate achieved by IA-SPA compared to the mean performance over 10 trials of random transmitter placement. The x-axis indicates the number of additional towers placed on top of the 9 existing AT&T towers in San Francisco, and the y-axis shows the resulting data rate.
Figure 14: Edge rate (5th percentile) comparison between IA-SPA and the mean performance over 10 trials of random placement. The x-axis indicates the number of additional towers placed on top of the 9 existing AT&T towers in San Francisco, and the y-axis shows the achieved data rate.
Figure 13 shows the average data rate as a function of the number of added transmitters. The blue curve corresponds to IA-SPA, while the orange curve shows the mean performance of 10 independent random placements. Even with the addition of a single transmitter, IA-SPA yields a substantial improvement over random placement.
Figure 14 presents the same comparison for the edge rate (5th percentile). IA-SPA again consistently outperforms random placement, with significant gains observed even when only a single transmitter is added.
TABLE V: Performance Metrics: Greedy vs. Random Placement (+5 added towers, San Francisco)
Metric Random IA-SPA Change
@@ IV-F San Francisco with fixed Initial Deployment
Mean Data Rate 30.56 40.29 $\uparrow$+31.82%
Std. Dev. 23.09 31.46 $\uparrow$+36.24%
Maximum Data Rate 156.56 176.17 $\uparrow$+12.52%
Edge Rate (5%-tile) 5.27 9.07 $\uparrow$+72.16%
@@
Mean Interference 8.93 7.92 $\downarrow$-11.29%
Std. Dev. 28.01 20.71 $\downarrow$-26.04%
Maximum Interference 1039.93 514.07 $\downarrow$-50.57%
Table V provides a detailed comparison for the case where 5 additional transmitters are deployed. IA-SPA improves the average data rate by approximately 30% relative to random placement (averaged over 10 trials), while the edge rate increases by 72%. Additionally, IA-SPA reduces interference, achieving an 11% decrease in mean interference compared to random placement. This illustrates that IA-SPA is capable of yielding promising performance in realistic continuous deployment scenarios.
@@ IV-G Key Takeaways from Results
The most significant results are found in the San Francisco AT&T scenario. Unlike other baselines that exhibit massive coverage gaps, the AT&T deployment is dense and well-structured. Despite this high-quality starting point, the IA-SPA achieved a 70% increase in mean throughput and a 88% improvement in the edge (5th percentile) data rate. These gains are particularly convincing because they were achieved while simultaneously slightly reducing the interference (by 2%). This proves that the algorithm is performing a high-precision optimization that disentangles overlapping signal footprints by exploiting the site-specific 3D geometry of the city.
The comparison between the two accumulation models highlights the distinction between raw signal capacity and interference-limited performance. While the $\mathcal{P}_{\text{SUM}}$ configuration provides an idealized upper bound for throughput by assuming perfect coordination, it lacks the implicit pressure to avoid overlapping coverage areas. This is reflected in the numerical results. While IA-SPA still achieves significant gains when using $\mathcal{P}_{\text{SUM}}$, the improvements are noticeably lower than those seen under $\mathcal{P}_{\text{MAX}}$ (e.g. 43% mean rate increase vs 70% for AT&T). This discrepancy suggests that by optimizing for aggregate power rather than strongest serving link, the algorithm may place transmitters in locations that provide redundant coverage, rather than expanding the network’s effective footprint. Consequently, the $\mathcal{P}_{\text{MAX}}$ objective not only provides a more realistic representation of signal behavior, but also serves as a more effective driver for identifying strategic sites that mitigate the specific geometric ”dead zone” characteristic for dense urban topologies.
The exclusionary zone and continuous deployment studies serve to demonstrate that the results of IA-SPA remain robust under various operational constraints. When considering an exclusionary zone, the algorithm achieved an 55% edge-rate improvement under placement constraints. Similarly we also showed that the algorithm is capable of incremental deployment by computing optimal transmitters to be added to an existing infrastructure, which resulted in a 29% improvement in data rate and a 30% reduction in interference.